OpenStudy (anonymous):

what is the set-builder notation of {2,3,5,7,11,13,17,19}

OpenStudy (zzr0ck3r):

\(\{p\in \mathbb{Z} \mid 2\le p \le 19 \text{ and } p \text{ is prime}\}\)

OpenStudy (zzr0ck3r):

there are many....

OpenStudy (zzr0ck3r):

\(\{p\in \mathbb{N} \mid p=2,3,5,7,11,13,17,19\}\)

OpenStudy (anonymous):

ohhh I was thinking that the difference btwn the first term is 1 the difference btwn the second and third term is 2 the difference btwn the third and 4th term is 3

OpenStudy (zzr0ck3r):

that wont work for the next one

OpenStudy (anonymous):

But not sure how I would apply the set builder notation

OpenStudy (zzr0ck3r):

13-11 = 2

OpenStudy (anonymous):

dang it u r right lol

OpenStudy (zzr0ck3r):

19-17 = 2

OpenStudy (zzr0ck3r):

if that worked you would be one of the richest men in the world

OpenStudy (anonymous):

hahahaha :P

OpenStudy (anonymous):

didnt even bother solving that far

OpenStudy (zzr0ck3r):

but for fun you should try and find the set builder notation for that one you described. ;)

OpenStudy (anonymous):

lol lemme sit on it

OpenStudy (anonymous):

Hey @zzr0ck3r I have the formula but how would you transform that into a set builder notation \[S_n=a_{n-1}+n\] Im gonna give it a shot... You can gear me into the right direction

OpenStudy (anonymous):

\[ \{ p\in {Z}\| ~p_=p_{n-1}+n\text{ where}~ 0\leq n\leq \infty \text{ and } p_0=2\}\] Or \[ \{ p\in {Z}\| ~p_n=p_{n-1}+n\text{ where}~ 0\leq n\leq \infty \text{ and } p_0=2\}\]

OpenStudy (zzr0ck3r):

:)